CHARACTERIZING TYPE I C*-ALGEBRAS VIA ENTROPY

NATHANIAL P. BROWN

Abstract. Let A be a separable unital C^{∗}-algebra. It is shown that A is type I if and only if the CNT-entropy of every inner automorphism of A is zero.